# LongCat-Flash-Omni Technical Report

### v1

@@ 4 Post-Training
The post-training stage is a critical phase that transforms the pre-trained foundation model into a task-adaptive, human-aligned system with strong instruction-following, multimodal reasoning, and interactive capabilities. While the pre-training phase enhances multimodal perception and world knowledge, post-training focuses on refining response alignment, controllability, and fidelity through a combination of supervised and preference-based optimization. This stage consists of two components: (1) Supervised Fine-Tuning (SFT), which equips the model with multimodal instruction-following, reasoning, and spoken interaction capabilities through high-quality and diverse instruction data, and (2) Reinforcement Learning (RL), which further enhances the model’s behavioral alignment, coherence, and consistency through Direct Preference Optimization (DPO) (Rafailov et al., 2023). Together, these procedures ensure that the final model not only demonstrates robust omni-modal understanding and reasoning capabilities but also generates responses that are semantically accurate, perceptually natural, and contextually coherent across both offline and online interaction scenarios.
@@ 4.1.2 Multimodal Interaction Data Curation
  • Model-Driven Automation To comprehensively cover diverse interaction scenarios, we first establish an ability taxonomy encompassing six major dimensions: memorization, understanding, analysis, creation, application, and entertainment. Guided by this taxonomy, we collect a mixture of public and in-house videos as the source data to ensure balanced coverage across abilities. For each video, we first perform scene segmentation using PySceneDetect (Breakthrough, 2025) to obtain a sequence of clips. For every clip, a powerful multimodal language model generates multiple rounds of progressively deep and context-aware QA pairs. In each round, subsequent queries build upon preceding conversations with logical progression, referential dependency, or anaphoric linkage to earlier turns. Such conversations encourage the model to reason over extended dialogue contexts, maintain entity and reference consistency in a natural, human-like audio-visual interaction setting. We then apply an automatic verification pipeline, employing an LLM-as-a-judge framework to evaluate and discard low-quality or inconsistent QA pairs. The remaining qualified samples are then composed into multi-turn dialogue, enabling the model to learn both intra-scene conversational continuity and inter-scene contextual transitions, thereby better reflecting real-world audio-visual interactions.
@@ 4.2 Reinforcement Learning
To further enhance the model’s multimodal capabilities and improve human alignment, we employ the Direct Preference Optimization (DPO) (Rafailov et al., 2023) during the reinforcement learning stage. The LongCat-Flash-Omni model supports multimodal inputs and enables parallel streaming outputs of both text and speech. However, most existing DPO variants are designed for text-only outputs or optimize text and speech separately. We argue that such decoupled optimization is suboptimal for maintaining coherence between textual and speech responses.
@@ 5.1 Multimodal Decoupling Framework
In multimodal scenarios, optimizing training performance is challenging due to the heterogeneity of both data and models. The challenge of data heterogeneity arises because the training data for LongCat-Flash-Omni includes speech, vision, and text, which show significant and dynamic differences in token distribution (Figure 7). Model heterogeneity is evident in the substantially different computational workloads across the three core components of LongCat-Flash-Omni: the vision encoder, the audio encoder, and the LLM decoder (Table 7). To address the aforementioned challenges, there are two primary optimization approaches. The first relies on Fully Sharded Data Parallelism (FSDP) (e.g., OrchMLLM (Zheng et al., 2025), veOmni (Ma et al., 2025)), which reduces static memory consumption through parameter sharding, avoids pipeline parallelism (PP) bubbles caused by model heterogeneity, and mitigates data heterogeneity via data-parallel (DP) group balancing. However, for LongCat-Flash-Omni, the total number of model parameters is too large for this approach to be practical.
The second approach is to decouple multimodal models across different parallel dimensions. For example, DistTrain (Zhang et al., 2025) mitigates model and data heterogeneity by separating model partitioning strategies and sorting data; PipeWeaver (Xue et al., 2025) reduces PP bubbles caused by heterogeneity through fine-grained data partitioning and dynamic pipeline scheduling; and Optimus (Feng et al., 2025) separates the parallelization strategies for encoders and LLMs, scheduling encoder computations into the LLM’s idle periods to eliminate bubbles caused by model heterogeneity. Building on the Optimus approach, we develop modality-decoupled parallelism (MDP), a simple yet effective multimodal training strategy. The core idea is to completely decouple the modality encoders and the LLM backbone at the distributed level, enabling independent scheduling and more efficient utilization of computational resources.
@@ 5.1.1 Modality-Decoupled Parallelism
In our implementation, we co-locate the modality encoders and the LLM decoder. The modality encoders utilize Hybrid Sharding Data Parallelism (HSDP)(Zhao et al., 2023) to reduce static memory, and full activation recomputation is employed to reduce activation memory usage. The LLM decoder adopts a combined distributed strategy including pipeline parallelism (PP), ZeRO-1 data parallelism (DP), context parallelism (CP), and expert parallelism (EP). To simplify data mapping between modality encoders and the LLM decoder, we introduce an InnerDP parallelism strategy, which further partitions modality data across all microbatches. The DP rank of modality encoders corresponds one-to-one with the DP rank of the LLM decoder, and the size of the InnerDP dimension is the product of the PP and CP of the LLM decoder (i.e., $d_{inner\_dp}=d_{lm\_cp}\times d_{lm\_pp}$, $d_{world\_size}=d_{dp}\times d_{inner\_dp}$). As shown in Figure 8, the MDP execution timeline consists of four phases:
@@ 5.3 Memory Optimization Strategies
We combine the techniques in Table 4: (i) a V-shaped PP schedule to bound concurrently alive activation micro-batches (Qi et al., 2024); (ii) selective recomputation for low-FLOP, high-activation operators such as SwiGLU and LayerNorm (DeepSeek-AI et al., 2025); (iii) memory-efficient permute, which moves aggregation of routing probabilities and hidden states into SwiGLU to reduce MoE unpermute memory (Shoeybi et al., 2019); (iv) fine-grained SM budgeting for NCCL communicators to reduce NCCL memory; and (v) Hybrid-Sharding Data Parallel (HSDP) for the modality encoder to further reduce its static footprint (Paszke et al., 2019). To accommodate potential EP load imbalance, we implement dynamic expert recomputation: when any rank is assigned too many tokens, we dynamically recompute its down-projection, freeing all MoE memory except the down-projection inputs on trigger. This prevents training crashes from expert imbalance while incurring only modest overhead.
@@ 7.1 Vision Capability Evaluation
In this section, we compare LongCat-Flash-Omni with omni-models (Gemini-2.5-Pro(Comanici et al., 2025), GPT-4o [^1] https://openai.com/index/hello-gpt-4o, Seed-1.6222https://seed.bytedance.com/en/seed1_6, and Qwen3-Omni(Xu et al., 2025)) and vision language models (Qwen3-VL(Yang et al., 2025) and Qwen2.5-VL-72B(Bai et al., 2025)) without thinking mode.
@@ 7.1.1 Image-to-Text Evaluation
  • General Domains: MMBench (Liu et al., 2024b), RealWorldQA (xAI, 2023), and MMStar (Chen et al., 2024b).
  • STEM & Reasoning: MathVista (Lu et al., 2024), MMMUval (Yue et al., 2024), and MMVet (Yu et al., 2024).
  • Multi-Image: BLINK (Fu et al., 2024a), MuirBench (Wang et al., 2024a), and Mantis (Jiang et al., 2024).
  • Text Recognition & Document Understanding: ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), and OCRBench (Liu et al., 2024c), OmniDocBenchEN/ZH (Ouyang et al., 2024).
  • Grounding & Counting: RefCOCO, RefCOCO+ (Kazemzadeh et al., 2014) (averaged in RefCOCO-avg), and CountBench(Paiss et al., 2023).
  • Graphical User Interface (GUI): VisualWebBench (Liu et al., 2024d), ScreenSpot-v2 (Wu et al., 2024a), and AndroidControl (Li et al., 2024a).
@@ 7.1.2 Video-to-Text Evaluation
  • Short Video: MVBench (Li et al., 2024b), NextQA (Xiao et al., 2021), and TempCompass (Liu et al., 2024e).
  • Long Video: VideoMME (Fu et al., 2024b) and LongVideoBench (Wu et al., 2024b).
  • STEM & Reasoning: MMVU (Zhao et al., 2025b), and Video-MMMU (Hu et al., 2025).
@@ 7.2.1 Base Model Evaluation
To systematically assess the audio capability of the base models produced in the pre-training stages (i.e., stage-1, stage-2, stage-3, and stage-4), we conduct evaluations with respect to the following three aspects: automatic speech recognition (ASR), text-to-speech (TTS), and speech continuation. For ASR evaluation, we employ the SPEECHIO_ASR_ZH00002 (speechio02)  [^3] https://github.com/SpeechColab/Leaderboard to test Chinese speech recognition, while utilizing the LibriSpeech (Panayotov et al., 2015) for English speech recognition assessment. The evaluation results are presented in Table 7. Remarkably, our model demonstrates competitive performance across these benchmark datasets, despite relying on discretized audio features.
@@
The TTS evaluation framework uses text samples from the speechio02 and LibriSpeech (Panayotov et al., 2015) test-clean/other datasets as input sources. The model uses discrete semantic-acoustic tokens as speech prompts, conducting speech token generation with text input as teacher-forcing guidance. The output speech tokens are reconstructed into waveform by the audio decoder. We compute the word error rate (WER) (for English) or character error rate (CER) (for Chinese) between transcription of the generated speech produced by a robust ASR system and the original input text as the measure for the evaluation. The results are shown in the right part of Table 7. We observe that the base models are capable of generating robust speech from textual input, providing a solid foundation for speech output in spoken and audio-visual interaction modes.
To evaluate speech continuation capabilities, we develope a specialized test set comprising $1,200$ synthesized speech samples derived from text data from the CMMLU (Li et al., 2023) benchmark. The assessment requires the model to generate appropriate responses to multiple-choice questions based on provided speech input. For text outputs, we directly measure response accuracy, while speech outputs undergo ASR transcription before accuracy evaluation. The evaluation protocol employs a 1-shot learning approach, with detailed results presented in Table 8. Our analysis reveals that the base models in different stage perform well in in-context speech continuation evaluation and there is only a negligible performance disparity between text and speech output.
@@ 7.2.2 Instruct Model Evaluation
For ASR evaluation, as presented in Table 9, we use LibriSpeech (Panayotov et al., 2015), AISHELL-1 (Bu et al., 2017), AISHELL-2 (Du et al., 2018), FLEURS (Conneau et al., 2023), CommonVoice15 (Ardila et al., 2019), and WenetSpeech (Zhang et al., 2022) as benchmarks. We observe that LongCat-Flash-Omni consistently achieves superior performance compared to other competitors, including Gemini-2.5-Pro (Comanici et al., 2025), GPT-4o-Audio, Qwen3-Omni-Instruct (Xu et al., 2025), Kimi-Audio (Ding et al., 2025), and Step-Audio-2-mini (Wu et al., 2025). For speech-to-text translation (S2TT) evaluation, we adopt CoVost2 (Wang et al., 2021). As shown in Table 9, LongCat-Flash-Omni exhibits strong S2TT capabilities among all models [^4] Kimi-audio defaults to ASR rather than executing the requested translation task.. Taken together, the speech recognition and translation results demonstrate that LongCat-Flash-Omni possesses robust and comprehensive fundamental speech understanding capabilities.
Audio Understanding As an omni-modal model, LongCat-Flash-Omni can effectively function as a native audio understanding model when vision input is not provided. We conduct a comprehensive evaluation of LongCat-Flash-Omni across a diverse set of audio comprehension tasks, including music, sound events, and speech understanding. Specifically, we use MMAU (Sakshi et al., 2024), VocalSound (Gong et al., 2022), TUT2017 (Mesaros et al., 2016), ClothoAQA (Lipping et al., 2022), Nonspeech7k (Rashid et al., 2023), CochlScene (Jeong and Park, 2022), and MELD (Poria et al., 2018) as benchmarks. The results, presented in Table 10, show that LongCat-Flash-Omni consistently outperforms most competing models across all evaluated dimensions, and achieves state-of-the-art performance in several benchmarks. These findings highlight that LongCat-Flash-Omni possesses advanced capabilities in understanding a wide spectrum of general and complex acoustic information, going well beyond conventional speech recognition.
Audio-to-Text Chat We evaluate the ability of LongCat-Flash-Omni to engage in text-based conversations driven by audio instructions using the OpenAudioBench (Li et al., 2025) and VoiceBench (Chen et al., 2024c) benchmarks. These benchmarks assess a wide range of capabilities, including world knowledge, domain-specific understanding, instruction following, and reasoning. To ensure standardized and reproducible results, we employ the official evaluation code provided by the benchmark authors. The results, presented in Table 11, show that LongCat-Flash-Omni achieves strong performance across all benchmark subsets, reaching state-of-the-art results in several cases. Overall, these results demonstrate LongCat-Flash-Omni ’s superior ability to conduct audio-driven conversations and perform complex reasoning tasks.

### v2

@@ 4 Post-Training
The post-training stage is a critical phase that transforms the pre-trained foundation model into a task-adaptive, human-aligned system with strong instruction-following, multimodal reasoning, and interactive capabilities. While the pre-training phase enhances multimodal perception and world knowledge, post-training focuses on refining response alignment, controllability, and fidelity through a combination of supervised and preference-based optimization. This stage consists of two components: (1) Supervised Fine-Tuning (SFT), which equips the model with multimodal instruction-following, reasoning, and spoken interaction capabilities through high-quality and diverse instruction data, and (2) Reinforcement Learning (RL), which further enhances the model’s behavioral alignment, coherence, and consistency through Direct Preference Optimization (DPO) (rafailov2023direct). Together, these procedures ensure that the final model not only demonstrates robust omni-modal understanding and reasoning capabilities but also generates responses that are semantically accurate, perceptually natural, and contextually coherent across both offline and online interaction scenarios.
@@ 4.1.2 Multimodal Interaction Data Curation
  • Model-Driven Automation To comprehensively cover diverse interaction scenarios, we first establish an ability taxonomy encompassing six major dimensions: memorization, understanding, analysis, creation, application, and entertainment. Guided by this taxonomy, we collect a mixture of public and in-house videos as the source data to ensure balanced coverage across abilities. For each video, we first perform scene segmentation using PySceneDetect (pyscenedetect) to obtain a sequence of clips. For every clip, a powerful multimodal language model generates multiple rounds of progressively deep and context-aware QA pairs. In each round, subsequent queries build upon preceding conversations with logical progression, referential dependency, or anaphoric linkage to earlier turns. Such conversations encourage the model to reason over extended dialogue contexts, maintain entity and reference consistency in a natural, human-like audio-visual interaction setting. We then apply an automatic verification pipeline, employing an LLM-as-a-judge framework to evaluate and discard low-quality or inconsistent QA pairs. The remaining qualified samples are then composed into multi-turn dialogue, enabling the model to learn both intra-scene conversational continuity and inter-scene contextual transitions, thereby better reflecting real-world audio-visual interactions.
@@ 4.2 Reinforcement Learning
To further enhance the model’s multimodal capabilities and improve human alignment, we employ the Direct Preference Optimization (DPO) (rafailov2023direct) during the reinforcement learning stage. The LongCat-Flash-Omni model supports multimodal inputs and enables parallel streaming outputs of both text and speech. However, most existing DPO variants are designed for text-only outputs or optimize text and speech separately. We argue that such decoupled optimization is suboptimal for maintaining coherence between textual and speech responses.
@@ 5.1 Multimodal Decoupling Framework
In multimodal scenarios, optimizing training performance is challenging due to the heterogeneity of both data and models. The challenge of data heterogeneity arises because the training data for LongCat-Flash-Omni includes speech, vision, and text, which show significant and dynamic differences in token distribution (Figure 7). Model heterogeneity is evident in the substantially different computational workloads across the three core components of LongCat-Flash-Omni: the vision encoder, the audio encoder, and the LLM decoder (Table 7). To address the aforementioned challenges, there are two primary optimization approaches. The first relies on Fully Sharded Data Parallelism (FSDP) (e.g., OrchMLLM (zheng2025orchestrate), veOmni (ma2025veomni)), which reduces static memory consumption through parameter sharding, avoids pipeline parallelism (PP) bubbles caused by model heterogeneity, and mitigates data heterogeneity via data-parallel (DP) group balancing. However, for LongCat-Flash-Omni, the total number of model parameters is too large for this approach to be practical.
The second approach is to decouple multimodal models across different parallel dimensions. For example, DistTrain (zhang2025disttrain) mitigates model and data heterogeneity by separating model partitioning strategies and sorting data; PipeWeaver (xue2025pipeweaver) reduces PP bubbles caused by heterogeneity through fine-grained data partitioning and dynamic pipeline scheduling; and Optimus (feng2025optimus) separates the parallelization strategies for encoders and LLMs, scheduling encoder computations into the LLM’s idle periods to eliminate bubbles caused by model heterogeneity. Building on the Optimus approach, we develop modality-decoupled parallelism (MDP), a simple yet effective multimodal training strategy. The core idea is to completely decouple the modality encoders and the LLM backbone at the distributed level, enabling independent scheduling and more efficient utilization of computational resources.
@@ 5.1.1 Modality-Decoupled Parallelism
In our implementation, we co-locate the modality encoders and the LLM decoder. The modality encoders utilize Hybrid Sharding Data Parallelism (HSDP)(zhao2023pytorchfsdpexperiencesscaling) to reduce static memory, and full activation recomputation is employed to reduce activation memory usage. The LLM decoder adopts a combined distributed strategy including pipeline parallelism (PP), ZeRO-1 data parallelism (DP), context parallelism (CP), and expert parallelism (EP). To simplify data mapping between modality encoders and the LLM decoder, we introduce an InnerDP parallelism strategy, which further partitions modality data across all microbatches. The DP rank of modality encoders corresponds one-to-one with the DP rank of the LLM decoder, and the size of the InnerDP dimension is the product of the PP and CP of the LLM decoder (i.e., $d_{inner\_dp}=d_{lm\_cp}\times d_{lm\_pp}$, $d_{world\_size}=d_{dp}\times d_{inner\_dp}$). As shown in Figure 8, the MDP execution timeline consists of four phases:
@@ 5.3 Memory Optimization Strategies
We combine the techniques in Table 4: (i) a V-shaped PP schedule to bound concurrently alive activation micro-batches (qi2024pipelineparallelismcontrollablememory); (ii) selective recomputation for low-FLOP, high-activation operators such as SwiGLU and LayerNorm (deepseekai2025deepseekv3technicalreport); (iii) memory-efficient permute, which moves aggregation of routing probabilities and hidden states into SwiGLU to reduce MoE unpermute memory (megatron-lm); (iv) fine-grained SM budgeting for NCCL communicators to reduce NCCL memory; and (v) Hybrid-Sharding Data Parallel (HSDP) for the modality encoder to further reduce its static footprint (paszke2019pytorch). To accommodate potential EP load imbalance, we implement dynamic expert recomputation: when any rank is assigned too many tokens, we dynamically recompute its down-projection, freeing all MoE memory except the down-projection inputs on trigger. This prevents training crashes from expert imbalance while incurring only modest overhead.
@@ 7.1 Vision Capability Evaluation
In this section, we compare LongCat-Flash-Omni with omni-models (Gemini-2.5-Pro(comanici2025gemini), GPT-4o [^1] https://openai.com/index/hello-gpt-4o, Seed-1.6222https://seed.bytedance.com/en/seed1_6, and Qwen3-Omni(xu2025qwen3)) and vision language models (Qwen3-VL(yang2025qwen3) and Qwen2.5-VL-72B(Qwen2.5-VL)) without thinking mode.
@@ 7.1.1 Image-to-Text Evaluation
  • General Domains: MMBench (liu2024mmbench), RealWorldQA (realworldqa), and MMStar (mmstar).
  • STEM & Reasoning: MathVista (lu2024mathvista), MMMUval (yue2023mmmu), and MMVet (yu2024mmvet).
  • Multi-Image: BLINK (fu2024blink), MuirBench (wang2024muirbench), and Mantis (jiang2024mantis).
  • Text Recognition & Document Understanding: ChartQA (masry2022chartqa), DocVQA (mathew2021docvqa), and OCRBench (liu2024ocrbench), OmniDocBenchEN/ZH (ouyang2024omnidocbench).
  • Grounding & Counting: RefCOCO, RefCOCO+ (kazemzadeh-2014-refcoco) (averaged in RefCOCO-avg), and CountBench(paiss2023countbench).
  • Graphical User Interface (GUI): VisualWebBench (liu2024visualwebbench), ScreenSpot-v2 (wu2024atlas), and AndroidControl (li2024androidcontrol).
@@ 7.1.2 Video-to-Text Evaluation
  • Short Video: MVBench (li2024mvbench), NextQA (xiao2021nextqa), and TempCompass (liu2024tempcompass).
  • Long Video: VideoMME (fu2024videomme) and LongVideoBench (wu2024longvideobench).
  • STEM & Reasoning: MMVU (zhao2025mmvu), and Video-MMMU (hu2025videommmu).
@@ 7.2.1 Base Model Evaluation
To systematically assess the audio capability of the base models produced in the pre-training stages (i.e., stage-1, stage-2, stage-3, and stage-4), we conduct evaluations with respect to the following three aspects: automatic speech recognition (ASR), text-to-speech (TTS), and speech continuation. For ASR evaluation, we employ the SPEECHIO_ASR_ZH00002 (speechio02)  [^3] https://github.com/SpeechColab/Leaderboard to test Chinese speech recognition, while utilizing the LibriSpeech (panayotov2015librispeech) for English speech recognition assessment. The evaluation results are presented in Table 7. Remarkably, our model demonstrates competitive performance across these benchmark datasets, despite relying on discretized audio features.
@@
The TTS evaluation framework uses text samples from the speechio02 and LibriSpeech (panayotov2015librispeech) test-clean/other datasets as input sources. The model uses discrete semantic-acoustic tokens as speech prompts, conducting speech token generation with text input as teacher-forcing guidance. The output speech tokens are reconstructed into waveform by the audio decoder. We compute the word error rate (WER) (for English) or character error rate (CER) (for Chinese) between transcription of the generated speech produced by a robust ASR system and the original input text as the measure for the evaluation. The results are shown in the right part of Table 7. We observe that the base models are capable of generating robust speech from textual input, providing a solid foundation for speech output in spoken and audio-visual interaction modes.
To evaluate speech continuation capabilities, we develope a specialized test set comprising $1,200$ synthesized speech samples derived from text data from the CMMLU (li2023cmmlu) benchmark. The assessment requires the model to generate appropriate responses to multiple-choice questions based on provided speech input. For text outputs, we directly measure response accuracy, while speech outputs undergo ASR transcription before accuracy evaluation. The evaluation protocol employs a 1-shot learning approach, with detailed results presented in Table 8. Our analysis reveals that the base models in different stage perform well in in-context speech continuation evaluation and there is only a negligible performance disparity between text and speech output.
@@ 7.2.2 Instruct Model Evaluation
For ASR evaluation, as presented in Table 9, we use LibriSpeech (panayotov2015librispeech), AISHELL-1 (bu2017aishell), AISHELL-2 (du2018aishell), FLEURS (conneau2023fleurs), CommonVoice15 (ardila2019common), and WenetSpeech (zhang2022wenetspeech) as benchmarks. We observe that LongCat-Flash-Omni consistently achieves superior performance compared to other competitors, including Gemini-2.5-Pro (comanici2025gemini), GPT-4o-Audio, Qwen3-Omni-Instruct (xu2025qwen3), Kimi-Audio (ding2025kimi), and Step-Audio-2-mini (wu2025step). For speech-to-text translation (S2TT) evaluation, we adopt CoVost2 (wang2021covost). As shown in Table 9, LongCat-Flash-Omni exhibits strong S2TT capabilities among all models [^4] Kimi-audio defaults to ASR rather than executing the requested translation task.. Taken together, the speech recognition and translation results demonstrate that LongCat-Flash-Omni possesses robust and comprehensive fundamental speech understanding capabilities.
Audio Understanding As an omni-modal model, LongCat-Flash-Omni can effectively function as a native audio understanding model when vision input is not provided. We conduct a comprehensive evaluation of LongCat-Flash-Omni across a diverse set of audio comprehension tasks, including music, sound events, and speech understanding. Specifically, we use MMAU (sakshi2024mmau), VocalSound (gong_vocalsound), TUT2017 (Mesaros2016_EUSIPCO), ClothoAQA (lipping2022clotho), Nonspeech7k (rashid2023nonspeech7k), CochlScene (jeong2022cochlscene), and MELD (poria2018meld) as benchmarks. The results, presented in Table 10, show that LongCat-Flash-Omni consistently outperforms most competing models across all evaluated dimensions, and achieves state-of-the-art performance in several benchmarks. These findings highlight that LongCat-Flash-Omni possesses advanced capabilities in understanding a wide spectrum of general and complex acoustic information, going well beyond conventional speech recognition.
Audio-to-Text Chat We evaluate the ability of LongCat-Flash-Omni to engage in text-based conversations driven by audio instructions using the OpenAudioBench (li2025baichuan) and VoiceBench (chen2024voicebench) benchmarks. These benchmarks assess a wide range of capabilities, including world knowledge, domain-specific understanding, instruction following, and reasoning. To ensure standardized and reproducible results, we employ the official evaluation code provided by the benchmark authors. The results, presented in Table 11, show that LongCat-Flash-Omni achieves strong performance across all benchmark subsets, reaching state-of-the-art results in several cases. Overall, these results demonstrate LongCat-Flash-Omni ’s superior ability to conduct audio-driven conversations and perform complex reasoning tasks.